OpenStudy (anonymous):

I NEED HELP.HOW DO I SOLVE THIS SOLVING BY SUBSTITUCION 5X+3Y=0 X-3Y=-18

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

substitution method = substitute... from your second eq: x=3y-18 substitute value of x in first eq: 5*(3y-18)+3y=0 solve it for y, then find x... let me know if there is a problem :)

OpenStudy (anonymous):

solve the second equation for x since it has a 1x as the number in front of the variable: x = 3y -18 plug the (3y - 18) in place of x in the first equation. (MAJOR RULE: You must always substitute into the other equation!!): 5(3y-18) +3y = 0 distribute: 15y - 90 + 3y =0 combine like terms and solve: 18y = 90 y = 5 Then plug in 5 for y in EITHER of the original equations. Choose the easier one: x - 3(5) =-18 then solve for x: x - 15 = -18 x = -3 Final answer: (-3, 5) you can check by plugging x and y into both equations to verify that it works for both. if it does, it has to be the right answer.